# Non-commutative Iwasawa theory of abelian varieties over global function fields

### v1

@@ 7. The order of vanishing of the characteristic element
In this section, we prove our main result on the order of vanishing of the characteristic element attached to the Selmer group over a $p$-adic Lie extension. We shall begin recalling the definition of these characteristic elements. As before, $G$ will denote a compact $p$-adic Lie group without $p$-torsion and $H$ is a closed normal subgroup of $G$ such that $\Gamma:=G/H\simeq{\mathbb{Z}}_{p}$. First, we introduce the characteristic element $\xi_{M}$ attached to $M\in\mathfrak{M}_{H}(G)$. In order to do this, we recall from [8] an equivalent definition of the category $\mathfrak{M}_{H}(G)$. A multiplicatively closed subset $\Sigma$ of $\Lambda(G)$ is said to be a left and right Ore set if, for each $f\in\Sigma$ and $x\in\Lambda(G)$, there exist $g_{1},g_{2}\in\Sigma$ and $y_{1},y_{2}\in\Lambda(G)$ such that
@@
[…]
It is shown in [8, Theorem 2.4] that the set $\Sigma$ is a multiplicatively closed left and right Ore set in $\Lambda(G)$. Since $p\notin\Sigma$, we also need to consider
@@
which is again a multiplicatively closed left and right Ore set in $\Lambda(G)$. We may therefore localise $\Lambda(G)$ at $\Sigma$ and at $\Sigma^{*}$ to obtain $\Lambda(G)_{\Sigma}$ and $\Lambda(G)_{\Sigma^{*}}=\Lambda(G)_{\Sigma}\left[\frac{1}{p}\right]$, respectively. It follows from [8, Proposition 2.3] that, given a finitely generated $\Lambda(G)$-module $M$, we have $M\in\mathfrak{M}_{H}(G)$ if and only if $M$ is $\Sigma^{*}$-torsion. Since the Iwasawa algebra $\Lambda(G)$ is noetherian and $G$ has no element of order $p$, $\Lambda(G)$ has finite global dimension equal to $\dim(G)+1$. Then $K_{0}(\Lambda(G))$ can be identified with the Grothendieck group of the category of all finitely generated $\Lambda(G)$-modules. It follows from the localisation sequence of $K$-theory for the Ore set $\Sigma^{*}$ that we have an exact sequence
@@
By [8, Proposition 3.4], the map $\partial_{G}$ is surjective. In view of this result, we can define the notion of a characteristic element for each $M$ in $\mathfrak{M}_{H}(G)$.
@@
Given $M$ in $\mathfrak{M}_{H}(G)$, a characteristic element for $M$ is any element $\xi_{M}\in K_{1}(\Lambda(G)_{\Sigma^{*}})$ satisfying
@@
We fix an algebraic closure ¯$\overline{{\mathbb{Q}}}_{p}$ of ${\mathbb{Q}}_{p}$. Let $\rho:G\to{\mathrm{GL}}_{n}(\mathcal{O})$ be an Artin representation, where $\mathcal{O}$ is the ring of integers of a finite extension of ${\mathbb{Q}}_{p}$ in ¯$\overline{{\mathbb{Q}}}_{p}$. We write $\Lambda_{\mathcal{O}}(\Gamma)$ for the Iwasawa algebra of $\Gamma$ with coefficients in $\mathcal{O}$, which we identify with the formal power series ring $\mathcal{O}[[T]]$ in an indeterminate $T$ by sending our fixed topological generator of $\Gamma$ to $1+T$. Let $\mathcal{Q}_{\mathcal{O}}(\Gamma)$ denote the field of fractions of $\Lambda_{\mathcal{O}}(\Gamma)$. If $M$ is a finitely generated $\Lambda(G)$-module, we put $M_{\mathcal{O}}=M\otimes_{{\mathbb{Z}}_{p}}\mathcal{O}$, and define
@@
endowed with the diagonal action of $G$, where $G$ acts on $\mathcal{O}^{n}$ via $\rho$.
@@
which maps $g$ to $\rho(g)\otimes\bar{g}$, where $\bar{g}$ denotes the image of $g$ in $\Gamma$. By [8, Lemma 3.3], this extends to a ring homomorphism
@@
From the functoriality of $K_{1}$-groups, this in turns induces a group homomorphism
@@
We then set $\Phi_{\rho}$ to be the composite map
(7.2) $\Phi_{\rho}:K_{1}(\Lambda(G)_{\Sigma^{*}})\to K_{1}(M_{n}(\mathcal{Q}_{\mathcal{O}}(\Gamma)))\stackrel{{\scriptstyle\text{Morita}}}{{\simeq}}K_{1}(\mathcal{Q}_{\mathcal{O}}(\Gamma))\simeq\mathcal{Q}_{\mathcal{O}}(\Gamma)^{\times}.$
Here, the first isomorphism is given by Morita’s equivalence. We also identify $\mathcal{Q}_{\mathcal{O}}(\Gamma)^{\times}$ with $\mathcal{Q}_{\mathcal{O}}(T)^{\times}$, where $\mathcal{Q}_{\mathcal{O}}(T)$ denotes the field of fractions of the power series ring $\mathcal{O}[[T]]$.
@@
Let $\xi\in K_{1}(\Lambda(G)_{\Sigma^{*}})$ and let $\rho$ be an Artin representation of $G$. Then there exists an integer $r_{\rho}(\xi)$ and $g(T)\in\mathcal{Q}_{\mathcal{O}}(T)^{\times}$ such that $g(0)\neq 0$ and
@@
We call $r_{\rho}(\xi)$ the order of vanishing of $\xi$ and denote by ${\mathrm{ord}}_{T=0}(\Phi_{\rho}(\xi))$.
We now relate the evaluation of the characteristic element of a module $M$ in the category $\mathfrak{M}_{H}(G)$ at Artin representations of $G$ to the Akashi series of the twists of $M$ by such representations defined by (7.1).
@@
where $\hat{\rho}$ denotes the contragradient representation of $\rho$, i.e. $\hat{\rho}(g)=\rho(g^{-1})^{t}$ for $g\in G$, where $t$ denotes the transpose matrix.
@@
By [8, Lemma 3.2], the module $\mathrm{tw}_{\hat{\rho}}(M)$ also lies in the category $\mathfrak{M}_{H}(G)$. Thus, $\mathrm{Ak}(\mathrm{tw}_{\hat{\rho}}(M))$ is well-defined, and the equality in the lemma is [8, (57)]. ∎
Given an Artin representation $\rho:G\to GL_{n}(\mathcal{O})$ of $G$ with coefficients in $\mathcal{O}$, let $F_{\rho}$ be the Galois extension of $F$ contained in $F_{\infty}$ such that ${\mathrm{Gal}}(F_{\infty}/F_{\rho})=\ker\rho$. We shall then write $G_{\rho}={\mathrm{Gal}}(F_{\infty}/F_{\rho})$. In the case the regular representation $\rho=\mathrm{reg}_{L}$ for a finite Galois extension $L$ of $F$ contained in $F_{\infty}$, we simply write $G_{L}={\mathrm{Gal}}(F_{\infty}/L)$ for $G_{\mathrm{reg}_{L}}$. We now introduce two hypotheses which will be useful for the computations in this section and in Section 9.
Hypothesis ($\mathbf{G}$). $H^{i}(G,A_{p^{\infty}}(F_{\infty}))$ is finite for all $i\geq 1$.
Hypothesis ($\mathbf{H}$). $H^{i}(H,A_{p^{\infty}}(F_{\infty}))$ is finite for all $i\geq 0$.
@@
Hypothesis $(\mathbf{G})$ is clearly satisfied if $A_{p^{\infty}}(F_{\infty})$ is finite. It also holds in the case when $F_{\infty}=F(A_{p^{\infty}})$ (see [2, Section 3]). Note that Hypothesis $(\mathbf{H})$ implies Hypothesis $(\mathbf{G})$.
Similarly, we write Hypothesis ($\mathbf{G_{\rho}}$) in case we replace $G$ with $G_{\rho}$, and Hypothesis $(\mathbf{G_{L}})$ in the case $\rho=\mathrm{reg}_{L}$ for a finite Galois extension $L$ of $F$ contained in $F_{\infty}$ with $G_{L}={\mathrm{Gal}}(F_{\infty}/L)$.
The next proposition provides a criterion for verifying Hypothesis $(\mathbf{H})$.
@@
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$. Suppose that $A_{p^{\infty}}(F^{\mathrm{cyc}})$ is finite and $H\simeq{\mathbb{Z}}_{p}^{r}$. Then Hypothesis $(\mathbf{H})$ holds, i.e., $H^{i}(H,A_{p^{\infty}}(F_{\infty}))$ is finite for all $i\geq 0$.
@@
The conclusion of the proposition is a consequence of the following algebraic observation: If $H\cong{\mathbb{Z}}_{p}^{r}$, then for a given $\Lambda(H)$-module $M$, one has $H_{i}(H,M)$ being finite for every $i\geq 1$ whenever $H_{0}(H,M)$ is finite. This latter assertion essentially follows from [33, P. 56-57], although the discussion given there is rather terse. We refer the reader to [22, Proof of Theorem 2.3] for a detailed proof of this fact. ∎
We recall that, by Remark 4.3, the finiteness condition on $A_{p^{\infty}}(F^{\mathrm{cyc}})$ automatically holds when $A$ is an elliptic curve without complex multiplication.
We now prove the following proposition which illustrates the significance of Hypothesis $(\mathbf{G})$.
@@
Let $A$ be an abelian variety defined over a global field of characteristic different from $p$. Suppose that Hypothesis $(\mathbf{G})$ holds. Then we have ${\mathrm{ord}}_{T=0}\left(f_{H^{i}(H,A_{p^{\infty}}(F_{\infty}))}\right)=0$ for all $i\geq 1$.
@@
To see that the characteristic series of $H^{i}(H,A_{p^{\infty}}(F_{\infty}))$ does not contribute a factor of $T$, it suffices to show that $H^{i}(H,A_{p^{\infty}}(F_{\infty}))^{\Gamma}$ is finite. Since $\mathrm{cd}_{p}\Gamma=1$, the spectral sequence
@@
degenerates yielding a short exact sequence
@@
for every $i\geq 1$. Since $H^{i}(G,A_{p^{\infty}}(F_{\infty}))$ is finite for every $i\geq 1$ by Hypothesis $(\mathbf{G})$, it follows that $H^{i}(H,A_{p^{\infty}}(F_{\infty}))^{\Gamma}$ is also finite for every $i\geq 1$, as required. ∎
Now, if $L$ is a finite Galois extension of $F$ contained in $F_{\infty}$, we write $\mathrm{reg}_{L}$ for the regular representation of ${\mathrm{Gal}}(L/F)$ and let $G_{L}$ denote the Galois group ${\mathrm{Gal}}(F_{\infty}/L)$.
@@
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Denote by $\xi_{A}=\xi_{X(A/F_{\infty})}$ the characteristic element of $X(A/F_{\infty})$. Suppose that Hypothesis $(\mathbf{G_{L}})$ holds. Then we have
@@
In addition, equality holds when Greenberg’s semi-simplicity conjecture 6.2 holds for $A/L$.
@@
Set $H_{L}=G_{L}\cap H$ and $\Gamma_{L}={\mathrm{Gal}}(L^{\mathrm{cyc}}/L)$. Fix an appropriate power of the generator of $\Gamma$ so that it is a generator of $\Gamma_{L}$. With these choices of generators, we have
@@
By the discussion in [21, Subsection 2.2 and Lemma 2.5], there is a restriction map
@@
such that $\mathrm{res}(\xi_{A})$ is a characteristic element of $X(A/F_{\infty})$ viewed in $\mathfrak{M}_{H_{L}}(G_{L})$. On the other hand, [21, Proposition 2.10] tells us that
@@
Hence, in view of the above observations, it suffices to consider the case $F=L$. It follows from Lemma 7.3 that
@@
for some $u(T)\in\Lambda(T)^{\times}$. Taking the hypothesis of the theorem into account, we may apply Proposition 7.6 to conclude that ${\mathrm{ord}}_{T=0}\left(f_{H^{j}(H,A_{p^{\infty}}(F_{\infty}))}\right)=0$ for all $j\geq 1$. Combining this latter observation into Theorem 5.5, we obtain
@@
By Proposition 6.3, the order of vanishing of $f_{{\mathrm{Sel}}(A/F^{\mathrm{cyc}})}$ at $T=0$ is greater or equal to ${\mathrm{corank}}_{{\mathbb{Z}}_{p}}({\mathrm{Sel}}(A/F))$, where equality holds if Conjecture 6.2 is valid. Hence it remains to show that the terms $f_{J_{v}(A/F^{\mathrm{cyc}})}$ at $v\in S^{\prime}$ does not contribute to the order of vanishing. For this, it suffices to show that $J_{v}(A/F^{\mathrm{cyc}})^{\Gamma}$ is finite. Indeed, noting that $\Gamma_{w}\simeq\Gamma\simeq{\mathbb{Z}}_{p}$ has $p$-cohomological dimension one, the inflation–restriction sequence gives us a surjection
@@
By Tate duality (cf. [26, Chapter I, Remark 3.6]), we see that $H^{1}(F_{v},A_{p^{\infty}})$ is dual to $A^{*}_{p^{\infty}}(F_{v})$, and the latter is a finite group by Mattuck’s theorem for local fields. This completes the proof of the theorem. ∎
We now present a corollary which relates the characteristic element to the $L$-function of $A$ in the intermediate subfields of $F_{\infty}/F$. For a finite extension $L$ of $F$ contained in $F_{\infty}$, we write $L_{S}(A/L,s)$ for the $L$-function of $A$ with Euler factors removed at the primes of $L$ lying above $S$. The following is our corollary.
@@
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Suppose that for a finite Galois extension $L$ of $F$ contained in $F_{\infty}$ the following statements are valid.
@@
    Hypothesis $(\mathbf{G_{L}})$ holds.
Then we have
@@
Under the finiteness of $\hbox{\cyr X}(A/L)$, we have ${\mathrm{corank}}_{{\mathbb{Z}}_{p}}({\mathrm{Sel}}(A/L))={\mathrm{rank}}_{\mathbb{Z}}A(L)$. On the other hand, under the finiteness of $\hbox{\cyr X}(A/L)$, the deep works of Artin–Tate [34], Milne [25] and Kato–Trihan [18] show that ${\mathrm{rank}}_{\mathbb{Z}}A(L)={\mathrm{ord}}_{s=1}L_{S}(A/L,s)$. The conclusion of the corollary then follows from combining these equalities with Theorem 7.7. ∎
It is now natural to ask whether similar results can be obtained for Artin representations which are not the regular representation. In order to do this, we introduce the following notation. Let $\rho:G\to GL_{n}(\mathcal{O})$ be an Artin representation of $G$ with coefficients in $\mathcal{O}$. Denote by $W_{\rho}$ a free $\mathcal{O}$-module of rank $n$ realizing $\rho$. Let $F_{\rho}$ be the Galois extension of $F$ contained in $F_{\infty}$ such that ${\mathrm{Gal}}(F_{\infty}/F_{\rho})=\ker\rho$. We shall then write $G_{\rho}={\mathrm{Gal}}(F_{\infty}/F_{\rho})$.
@@
Let $A$ be an abelian variety defined over a global field of characteristic different from $p$. Suppose that Hypothesis $(\mathbf{G_{\rho}})$ holds. Then we have ${\mathrm{ord}}_{T=0}\left(f_{H^{i}(H,\mathrm{tw}_{\rho}(A_{p^{\infty}})(F_{\infty}))}\right)=0$ for all $i\geq 1$.
@@
Let $H_{\rho}=G_{\rho}\cap H$. Consider the following spectral sequence
@@
Since $H/H_{\rho}$ is finite, it remains to show that the characteristic series of $H^{i}(H_{\rho},\mathrm{tw}_{\rho}(A_{p^{\infty}})(F_{\infty}))$ does not contribute a factor of $T$. By the definition of $H_{\rho}$, the group $H_{\rho}$ acts trivially on $W_{\rho}$, and so we have

### v2

@@ 7. The order of vanishing of the characteristic element
In this section, we prove our main result concerning the order of vanishing of the characteristic element associated with the Selmer group over a $p$-adic Lie extension. We begin by recalling the definition of these characteristic elements.
As before, let $G$ denote a compact $p$-adic Lie group without $p$-torsion, and let $H$ be a closed normal subgroup of $G$ such that $\Gamma:=G/H\simeq{\mathbb{Z}}_{p}$. First, we introduce the characteristic element $\xi_{M}$ attached to $M\in\mathfrak{M}_{H}(G)$. To this end, we recall from [9] an equivalent definition of the category $\mathfrak{M}_{H}(G)$. A multiplicatively closed subset $\Sigma\subset\Lambda(G)$ is said to be a left and right Ore set if, for each $f\in\Sigma$ and $x\in\Lambda(G)$, there exist $g_{1},g_{2}\in\Sigma$ and $y_{1},y_{2}\in\Lambda(G)$ such that
@@
$\Sigma=\{f\in\Lambda(G):\Lambda(G)/\Lambda(G)f\text{ is a finitely generated $\Lambda(H)$-module}\}.$
It is shown in [9, Theorem 2.4] that $\Sigma$ is a multiplicatively closed left and right Ore set in $\Lambda(G)$. Since $p\notin\Sigma$, we also consider the enlarged set
@@
which is again a multiplicatively closed left and right Ore set in $\Lambda(G)$. We may therefore localise $\Lambda(G)$ at $\Sigma$ and at $\Sigma^{*}$ to obtain $\Lambda(G)_{\Sigma}$ and $\Lambda(G)_{\Sigma^{*}}=\Lambda(G)_{\Sigma}\left[\frac{1}{p}\right]$, respectively. It follows from [9, Proposition 2.3] that a finitely generated $\Lambda(G)$-module $M$ lies in $\mathfrak{M}_{H}(G)$ if and only if $M$ is $\Sigma^{*}$-torsion. Since the Iwasawa algebra $\Lambda(G)$ is Noetherian and $G$ has no element of order $p$, the global dimension of $\Lambda(G)$ is finite. From [9], $K_{0}(\Lambda(G))$ can be identified with the Grothendieck group of the category of all finitely generated $\Lambda(G)$-modules. From the localisation sequence in $K$-theory for the Ore set $\Sigma^{*}$, we obtain the following exact sequence
@@
By [9, Proposition 3.4], the map $\partial_{G}$ is surjective. In light of this result, we can define a characteristic element for each $M\in\mathfrak{M}_{H}(G)$.
@@
Given $M\in\mathfrak{M}_{H}(G)$, a characteristic element for $M$ is any element $\xi_{M}\in K_{1}(\Lambda(G)_{\Sigma^{*}})$ satisfying
@@
We fix an algebraic closure $\overline{{\mathbb{Q}}}_{p}$ of ${\mathbb{Q}}_{p}$. Let
$\rho:G\to{\mathrm{GL}}_{n}(\mathcal{O})$
be an Artin representation, where $\mathcal{O}$ is the ring of integers of a finite extension of ${\mathbb{Q}}_{p}$ in $\overline{{\mathbb{Q}}}_{p}$. We write $\Lambda_{\mathcal{O}}(\Gamma)$ for the Iwasawa algebra of $\Gamma$ with coefficients in $\mathcal{O}$, which we identify with the formal power series ring $\mathcal{O}[[T]]$ by sending a fixed topological generator of $\Gamma$ to $1+T$. Let $\mathcal{Q}_{\mathcal{O}}(\Gamma)$ denote the field of fractions of $\Lambda_{\mathcal{O}}(\Gamma)$. If $M$ is a finitely generated $\Lambda(G)$-module, we define $M_{\mathcal{O}}=M\otimes_{{\mathbb{Z}}_{p}}\mathcal{O}$, and set
@@
where ${\mathcal{O}}^{n}$ is endowed with the action of $G$ via $\rho$, and $G$ acts diagonally on the tensor product.
@@
which maps each $g$ to $\rho(g)\otimes\bar{g}$, where $\bar{g}$ denotes the image of $g$ in $\Gamma$. By [9, Lemma 3.3], this extends to a ring homomorphism
@@
By the functoriality of $K_{1}$-groups, this in turn induces a group homomorphism
@@
We define $\Phi_{\rho}$ to be the composite map
(7.2) $\Phi_{\rho}:K_{1}(\Lambda(G)_{\Sigma^{*}})\to K_{1}(M_{n}(\mathcal{Q}_{\mathcal{O}}(\Gamma)))\stackrel{{\scriptstyle\text{Morita}}}{{\simeq}}K_{1}(\mathcal{Q}_{\mathcal{O}}(\Gamma))\simeq\mathcal{Q}_{\mathcal{O}}(\Gamma)^{\times},$
where the first isomorphism is given by Morita’s equivalence. We also identify $\mathcal{Q}_{\mathcal{O}}(\Gamma)^{\times}$ with $\mathcal{Q}_{\mathcal{O}}(T)^{\times}$, where $\mathcal{Q}_{\mathcal{O}}(T)$ denotes the field of fractions of $\mathcal{O}[[T]]$.
@@
Let $\xi\in K_{1}(\Lambda(G)_{\Sigma^{*}})$, and let $\rho$ be an Artin representation of $G$. Then there exists an integer $r_{\rho}(\xi)$ and $g(T)\in\mathcal{Q}_{\mathcal{O}}(T)^{\times}$, with $g(0)\neq 0$, such that
@@
We call $r_{\rho}(\xi)$ the order of vanishing of $\xi$ at $\rho$, and denote it by ${\mathrm{ord}}_{T=0}(\Phi_{\rho}(\xi))$.
We now relate the evaluation of the characteristic element of a module $M\in\mathfrak{M}_{H}(G)$ at Artin representations of $G$ to the Akashi series of the twists of $M$ by such representations defined by (7.1).
@@
where $\hat{\rho}$ denotes the contragradient representation of $\rho$, defined by $\hat{\rho}(g)=\rho(g^{-1})^{t}$ for $g\in G$, where $t$ denotes the transpose matrix.
@@
By [9, Lemma 3.2], the twisted module $\mathrm{tw}_{\hat{\rho}}(M)$ also lies in the category $\mathfrak{M}_{H}(G)$. Hence, the Akashi series $\mathrm{Ak}(\mathrm{tw}_{\hat{\rho}}(M))$ is well-defined, and the claimed equality follows from [9, (57)]. ∎
Given an Artin representation $\rho:G\to GL_{n}(\mathcal{O})$ of $G$ with coefficients in $\mathcal{O}$, let $F_{\rho}$ be the Galois extension of $F$ contained in $F_{\infty}$ such that ${\mathrm{Gal}}(F_{\infty}/F_{\rho})=\ker\rho$. We then write $G_{\rho}={\mathrm{Gal}}(F_{\infty}/F_{\rho})$. In the case where $\rho=\mathrm{reg}_{L}$ is the regular representation corresponding to a finite Galois extension $L$ of $F$ contained in $F_{\infty}$, we simply write $G_{L}={\mathrm{Gal}}(F_{\infty}/L)$ for $G_{\mathrm{reg}_{L}}$. We now introduce two assumptions ($\mathbf{G}$) and ($\mathbf{H}$), which will be useful for the computations in this section and in Section 9.
$\mathbf{(G)}:\quad H^{i}(G,A_{p^{\infty}}(F_{\infty}))\text{ is finite for all }i\geq 1.$
$\mathbf{(H)}:\quad H^{i}(H,A_{p^{\infty}}(F_{\infty}))\text{ is finite for all }i\geq 0.$
@@
$(\mathbf{G})$ is clearly satisfied if $A_{p^{\infty}}(F_{\infty})$ is finite. It also holds in the case when $F_{\infty}=F(A_{p^{\infty}})$ (see [2, Section 3]). Note that $(\mathbf{H})$ implies $(\mathbf{G})$.
Similarly, we write ($\mathbf{G_{\rho}}$) when replacing $G$ with $G_{\rho}$, and $(\mathbf{G_{L}})$ in the case where $\rho=\mathrm{reg}_{L}$ for a finite Galois extension $L$ of $F$ contained in $F_{\infty}$, with $G_{L}={\mathrm{Gal}}(F_{\infty}/L)$.
The following proposition provides a criterion for verifying assumption $(\mathbf{H})$.
@@
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$. Suppose that $A_{p^{\infty}}(F^{\mathrm{cyc}})$ is finite and $H\simeq{\mathbb{Z}}_{p}^{r}$. Then $(\mathbf{H})$ holds, i.e., $H^{i}(H,A_{p^{\infty}}(F_{\infty}))$ is finite for all $i\geq 0$.
@@
The conclusion follows from the following algebraic observation: If $H\cong{\mathbb{Z}}_{p}^{r}$, then for a any $\Lambda(H)$-module $M$, the cohomology groups $H_{i}(H,M)$ are finite for all $i\geq 1$, provided that $H_{0}(H,M)$ is finite. This result essentially follows from [41, pp. 56–57], although the argument there is somewhat terse. For a detailed proof, we refer the reader to [26, Proof of Theorem 2.3]. ∎
We also recall, by Remark 4.3, that the finiteness condition on $A_{p^{\infty}}(F^{\mathrm{cyc}})$ automatically holds when $A$ is an elliptic curve without complex multiplication.
We now prove the following proposition, which highlights the significance of assumption $(\mathbf{G})$.
@@
Let $A$ be an abelian variety defined over a global field of characteristic different from $p$. Suppose that condition $(\mathbf{G})$ holds. Then, for all $i\geq 1$, we have ${\mathrm{ord}}_{T=0}\left(f_{H^{i}(H,A_{p^{\infty}}(F_{\infty}))}\right)=0$.
@@
To show that the characteristic series of $H^{i}(H,A_{p^{\infty}}(F_{\infty}))$ does not contribute a factor of $T$, it suffices to show that $H^{i}(H,A_{p^{\infty}}(F_{\infty}))^{\Gamma}$ is finite for all $i\geq 1$. Since $\mathrm{cd}_{p}\Gamma=1$, the Hochschild-Serre spectral sequence
@@
degenerates to a short exact sequence
@@
for every $i\geq 1$. Since $H^{i}(G,A_{p^{\infty}}(F_{\infty}))$ is finite for every $i\geq 1$ by assumption $(\mathbf{G})$, it follows that $H^{i}(H,A_{p^{\infty}}(F_{\infty}))^{\Gamma}$ is also finite for every $i\geq 1$, as required. ∎
Now, if $L$ is a finite Galois extension of $F$ contained in $F_{\infty}$, we write $\mathrm{reg}_{L}$ for the regular representation of ${\mathrm{Gal}}(L/F)$, and let $G_{L}={\mathrm{Gal}}(F_{\infty}/L)$.
@@
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Denote by $\xi_{A}=\xi_{X(A/F_{\infty})}$ the characteristic element of $X(A/F_{\infty})$. Suppose that the condition $(\mathbf{G_{L}})$ holds. Then we have
@@
Moreover, equality holds if Greenberg’s semi-simplicity conjecture 6.2 is valid for $A/L$.
@@
Set $H_{L}=G_{L}\cap H$ and $\Gamma_{L}={\mathrm{Gal}}(L^{\mathrm{cyc}}/L)$. Fix a suitable power of the chosen topological generator of $\Gamma$ so that it is a generator of $\Gamma_{L}$. With this choice, we have
@@
As discussed in [27, Subsection 2.2 and Lemma 2.5], there exists a restriction map
@@
such that $\mathrm{res}(\xi_{A})$ is a characteristic element of $X(A/F_{\infty})$ viewed in $\mathfrak{M}_{H_{L}}(G_{L})$. Furthermore, [27, Proposition 2.10] gives the identity
@@
Thus, by the observations above, it suffices to treat the case $F=L$. By Lemma 7.3, we have
@@
for some $u(T)\in\Lambda(T)^{\times}$. Under the hypotheses of the theorem, we apply Proposition 7.6 to conclude that ${\mathrm{ord}}_{T=0}\left(f_{H^{j}(H,A_{p^{\infty}}(F_{\infty}))}\right)=0$ for all $j\geq 1$. Combining this with Theorem 5.5, we obtain
@@
By Proposition 6.3, we know that the order of vanishing of $f_{{\mathrm{Sel}}(A/F^{\mathrm{cyc}})}$ at $T=0$ is at least ${\mathrm{corank}}_{{\mathbb{Z}}_{p}}({\mathrm{Sel}}(A/F))$, where equality holds if Conjecture 6.2 is valid. It therefore remains to show that the terms $f_{J_{v}(A/F^{\mathrm{cyc}})}$ for $v\in S^{\prime}$ do not contribute to the order of vanishing. To see this, it suffices to show that $J_{v}(A/F^{\mathrm{cyc}})^{\Gamma}$ is finite. Indeed, since $\Gamma_{w}\simeq\Gamma\simeq{\mathbb{Z}}_{p}$ has $p$-cohomological dimension one, the inflation–restriction sequence yields a surjection
@@
By Tate local duality (cf. [33, Chapter I, Remark 3.6]), $H^{1}(F_{v},A_{p^{\infty}})$ is dual to $A^{*}_{p^{\infty}}(F_{v})$, which is finite by [5, Proposition 3.2]. This completes the proof. ∎
We now present a corollary that relates the characteristic element to the $L$-function of $A$ over the intermediate subfields of $F_{\infty}/F$. For a finite extension $L$ of $F$ contained in $F_{\infty}$, we write $L_{S}(A/L,s)$ for the $L$-function of $A$ with Euler factors at the primes of $L$ lying above $S$ removed. The following is our corollary.
@@
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Suppose that for a finite Galois extension $L$ of $F$ contained in $F_{\infty}$, the following conditions hold:
@@
    Condition $(\mathbf{G_{L}})$ holds.
Then
@@
Under the finiteness of $\hbox{\cyr X}(A/L)$, we have
${\mathrm{corank}}_{{\mathbb{Z}}_{p}}({\mathrm{Sel}}(A/L))={\mathrm{rank}}_{\mathbb{Z}}A(L).$
Moreover, the deep results of Artin–Tate [42], Milne [32], and Kato–Trihan [22] imply that
${\mathrm{rank}}_{\mathbb{Z}}A(L)={\mathrm{ord}}_{s=1}L_{S}(A/L,s),$
provided that $\hbox{\cyr X}(A/L)$ is finite. The conclusion then follows from combining these equalities with Theorem 7.7. ∎
It is now natural to ask whether similar results hold for non-regular Artin representations. To explore this, we introduce the following notation. Let $\rho:G\to GL_{n}(\mathcal{O})$ be an Artin representation of $G$ with coefficients in $\mathcal{O}$. Let $W_{\rho}$ be a free $\mathcal{O}$-module of rank $n$ realising $\rho$. Denote by $F_{\rho}$ the Galois extension of $F$ contained in $F_{\infty}$ such that ${\mathrm{Gal}}(F_{\infty}/F_{\rho})=\ker\rho$. We then write $G_{\rho}={\mathrm{Gal}}(F_{\infty}/F_{\rho})$.
@@
Let $A$ be an abelian variety defined over a global field of characteristic different from $p$. Suppose that condition $(\mathbf{G_{\rho}})$ holds. Then, for all $i\geq 1$, we have
${\mathrm{ord}}_{T=0}\left(f_{H^{i}(H,\mathrm{tw}_{\rho}(A_{p^{\infty}})(F_{\infty}))}\right)=0.$
@@
Let $H_{\rho}=G_{\rho}\cap H$. Consider the spectral sequence
@@
Since $H/H_{\rho}$ is finite, it suffices to show that the characteristic series of $H^{i}(H_{\rho},\mathrm{tw}_{\rho}(A_{p^{\infty}})(F_{\infty}))$ does not contribute a factor of $T$. By the definition of $H_{\rho}$, this group acts trivially on $W_{\rho}$, and so we have